# STEP: A Unified Spiking Transformer Evaluation Platform for Fair and Reproducible Benchmarking

### v1

@@ 1 Introduction
Spiking Neural Networks (SNNs) are a biologically inspired paradigm that simulate neural information processing via discrete spikes. These networks excel not only at static image tasks but also in modeling dynamic and temporally structured data [1]. Their event-driven nature contributes to high energy efficiency and strong biological plausibility. However, applying SNNs to deep learning architectures—particularly Transformers—remains challenging due to their non-differentiability, limited scalability, and training instability.
@@
In parallel, Artificial Neural Networks (ANNs) have seen tremendous advances through architectural innovations. ResNet [2] introduced residual learning to ease optimization in deep networks, while Recurrent Neural Networks (RNNs) captured sequential dependencies. The Transformer architecture [3] unified these advances by leveraging self-attention, enabling efficient parallel modeling of long-range dependencies. Vision Transformer (ViT) [4] further demonstrated the potential of attention mechanisms in visual tasks. Drawing inspiration from these architectures, the SNN community has proposed Spiking ResNet [5], SEW-ResNet [6], and spiking RNN variants [7, 8]. Recently, attention-based spiking models such as Spikformer [9], QKFormer [10], and SpikingResformer [11] have emerged. The Spike-Driven Transformer series [12, 13, 14] improves both efficiency and scalability, enabling applications in image segmentation and object detection.
@@
Despite advancements, several key challenges persist in Spiking Transformers (STs). First, the performance gap between STs and traditional ANNs remains unclear, especially regarding their unique advantages on temporal or relatively complicated data. A systematic evaluation across diverse datasets—static (e.g., ImageNet), event-driven (e.g., DVS-CIFAR10), and sequential (e.g., SCIFAR10)—is essential for assessing their potential. Second, STs consist of multiple interacting components, including spike encoders, neuron models, surrogate gradients, attention modules, and MLP heads, yet the contribution of each module is underexplored. Module-wise ablation is critical for understanding trade-offs and guiding optimization. Third, while SNNs inherently offer energy benefits through sparse, binary spike-based computation, direct comparisons to quantized Transformers are scarce. Quantifying the energy-performance trade-off is necessary to assess the practical utility of spiking models. Moreover, inconsistencies across development frameworks, such as SpikingJelly [15], BrainCog [16], and BrainPy [17], further hinder progress by complicating reproducibility, hyperparameter tuning, and fair model comparison. Currently, no unified platform exists for evaluating Spiking Transformers across tasks like classification, segmentation, and detection.
@@
To address these challenges, we introduce the Spiking Transformer Evaluation Platform (STEP), a unified benchmarking framework for building, evaluating, and comparing Spiking Transformers. STEP integrates representative implementations, supports modular component replacement, and enables consistent evaluation across visual tasks. It provides both training-from-scratch and pretraining–finetuning pipelines, and supports integration with backends such as SpikingJelly, BrainCog, and BrainPy. Moreover, leveraging MMSegmentation [18] and MMDetection [19], STEP extends support to dense prediction tasks. Our main contributions are as follows:
@@ 2 Preliminary
To enable spike-based processing, input signals are transformed into temporal spike trains via encoding schemes such as direct, rate, time-to-first-spike (TTFS), and phase encoding [20, 21, 22]. A detailed overview of encoding methods is provided in Appendix A.1.
@@
Spiking neurons transmit information via discrete spikes triggered by membrane potential dynamics. The Leaky Integrate-and-Fire (LIF) model [23] is widely used due to its simplicity and biological plausibility:
$V[t]=V[t-1]+\frac{1}{\tau}(X[t]-V[t-1]),\quad\text{if }V[t]\geq V_{th},\text{ emit spike and reset.}$. (1)
Variants like PLIF [24] and GLIF [25] enhance adaptability with learnable decay or gated mechanisms. Further details are provided in Appendix A.2.
@@
Recent ST models propose lightweight attention [12, 13], hierarchical designs (e.g., QKFormer [10]), and multi-task heads (e.g., FCN [26], FPN [27]) to enhance performance across modalities. These improvements drive STs toward practical deployment while retaining neuromorphic efficiency.
@@ 3.2 Broad Dataset Compatibility
Our benchmark supports a wide spectrum of datasets, encompassing static (e.g., ImageNet [28], CIFAR10/100 [29]), event-based (e.g., DVS-CIFAR10 [30], N-Caltech101 [31]), and sequential inputs. It also integrates sequential classification tasks to assess temporal modeling capabilities. For dense prediction tasks, we provide plug-and-play support for SNN-adapted segmentation and detection models, such as FCN [26] and FPN [27], based on MMSeg and MMDet toolchains.
@@ 3.3 Multi-task Adaptation
While early Spiking Transformers (e.g., Spikformer [9], TIM [32]) were largely limited to classification, recent efforts such as Spike-Driven Transformer V2 have expanded their scope to include dense vision tasks. Our benchmark extends this trajectory by enabling flexible configuration across classification, segmentation, and detection pipelines within a unified framework.
@@ 3.4 Backend-Agnostic Integration
To maximize accessibility and reuse, the benchmark supports multiple backends including SpikingJelly [15], BrainCog [16], and BrainPy, among others. This backend-agnostic design ensures compatibility with different development environments and facilitates cross-framework reproducibility. Together, these components form a robust, extensible, and task-agnostic benchmarking framework. By supporting flexible model construction, diverse datasets, multi-task adaptation, and backend interoperability, our benchmark provides a solid foundation for future development and systematic evaluation of Spiking Transformer architectures.
@@ 4.1 Reproduction
Spikformer [9] 128 4 400 95.12 (95.41) 77.37 (78.21)
SDT [12] 128 4 400 95.77 (95.60) 78.29 (78.40)
QKFormer [10] 128 4 400 96.24 (96.18) 79.72 (81.15)
Spikingformer [33] 128 4 400 95.53 (95.81) 79.12 (79.21)
Spikformer + SEMM [34] 128 4 400 94.98 (95.78) 77.59 (79.04)
Spiking Wavelet [35] 128 4 400 95.31 (96.10) 76.99 (79.30)
SGLFormer [36] 16 4 400 95.88 (96.76) 80.61 (82.26)
SpikingResformer [33]∗∗ 128 4 400 95.69 (97.40) 79.45 (85.98)
Tab. 1 presents our reproduced results on CIFAR-10 and CIFAR-100 [29]. All models are trained using the same optimizer, learning rate, batch size (unless otherwise constrained), training epochs, and random seed. Experiments are conducted on NVIDIA A100 GPUs with 40GB memory.
Overall, our reproduced results are consistent with the original papers. Some models, such as QKFormer, even outperform their reported results, suggesting strong reproducibility. Discrepancies stem mainly from (i) implementation differences, e.g., SpikingResformer originally uses transfer learning, while our setup employs end-to-end training; and (ii) memory limitations, e.g., SGLFormer requires a smaller batch size. To ensure fairness, we avoid dataset- or model-specific tuning and apply a uniform experimental protocol across all baselines.
@@ 4.2.1 Classification:  ImageNet-1K
For ImageNet-1K [28] we evaluate only Spikformer [34] and QKFormer [10]: the former is the seminal Spiking-Transformer baseline, while the latter introduces a hierarchical pyramid and currently delivers SOTA accuracy among SNN-based Transformers. Concentrating our limited GPU budget on these two “end-points” lets us cover the full architectural spectrum without incurring the prohibitive cost of training several similar intermediate models. Because ImageNet-1K is orders of magnitude larger and more complex than CIFAR-10/100—and because Spikformer and QKFormer differ greatly in parameter count and memory footprint—forcing a single batch size and epoch schedule would either overflow A100 GPU memory or demand untenable compute. We therefore keep each model’s published regime (QKFormer: 200 epochs × 32/GPU; Spikformer: 300 epochs × 24/GPU), while unifying every other hyper-parameter under a single script; the differing batch sizes and epoch counts are thus an intentional, resource-aware decision rather than an oversight.
@@ 4.2.2 Segmentation: ADE20K
QKFormer [10] 256 4 200 73.88 - - -
Spikformer [33] 192 4 300 73.69 69.80 23.51 31.43
Spikformer + SEMM [34] - - - - 63.41 13.13 19.76
SDT [12] - - - - 63.45 12.08 17.17
For semantic segmentation on ADE20K [37], only SDT [12] had been previously evaluated. We conduct a fair comparison by retraining Spikformer  [34]and SEMM [38] without pretraining. Interestingly, both Spikformer variants outperform SDT under identical settings, despite SDT’s original paper reporting strong performance with pretraining. This implies that with appropriate initialization strategies, Spikformer-based models could potentially surpass existing baselines in dense prediction tasks. The segmentation result can be viewed in Fig. 3
@@ 5.1 Neuron Model Evaluation
To investigate the impact of different spiking neuron types on model performance, we replace the default LIF neuron with five widely used variants: PLIF [24], CLIF [39], GLIF [25], and KLIF [40]. These models extend the basic LIF neuron [23] by incorporating enhancements such as learnable time constants, gating mechanisms, and surrogate gradient improvements. To quantify the influence of neuron choice, we replace the default LIF cell with four mainstream variants—PLIF [24], CLIF [39], GLIF [25], and KLIF [40]. Each variant augments the canonical LIF formulation [23] with additional biological or optimization benefits, ranging from a learnable membrane constant (PLIF) to gated internal states (GLIF) and surrogate-gradient refinements (CLIF, KLIF). Detailed explaination can be found in Appendix A.2.
@@
Spikformer [9] 95.12 95.38 95.41 95.85 96.06
SDT [12] 95.77 95.49 95.45 95.63 95.91
Spikformer + SEMM [34] 94.98 95.44 95.78 95.59 95.66
@@ 5.2 Sequence Modeling
Recent studies [41] question the ability of standard Transformers to model long-range temporal dependencies, prompting alternatives like Spiking SSM [42]. Datasets such as sCIFAR [43] and (p)sMNIST [44] serialize 2D images into 1D sequences, emphasizing temporal structure. Spiking SSM processes inputs at the pixel level (e.g., 784 steps for a full MNIST image), incurring high computational costs. To adapt Spiking Transformers for serialized inputs, we replace 2D convolutions in the SPS module with 1D convolutions.
@@
FlexTCN [45] No 99.62 98.63 80.82
SMPConv [46] No 99.75 99.10 84.86
LMUformer [47] No - 98.55 -
ViT [4] No 98.00 97.73 74.95
ViT[4] + SPS ∗∗ No 99.19 98.19 85.62
SpikingSSM [42] Yes 99.60 98.40 -
SpikingLMUformer [47] Yes - 97.92 -
Spikformer [9] Yes 98.84 97.97 84.26
SDT [12] Yes 98.77 97.80 82.31
Spikformer + SEMM [34] Yes 99.33 98.46 85.61
@@ 5.3 Encoding Schemes
RGB inputs can be converted to spikes using four common encoding schemes: direct, phase, rate, and TTFS [20, 21, 22]. As shown in Tab. LABEL:tab:encoding_test, direct encoding, a lossless approach ensures that no information is lost—which repeats the entire image at every time step matches current Spiking Transformers that compute attention independently across steps, and therefore yields the highest Top-1 accuracy on all models. In contrast, the sparser phase, rate, and TTFS encodings reduce spike density and weaken spatial coherence, leading to lower accuracy and highlighting the need for future architectures with temporally aware attention or recurrent mechanisms.
@@ 5.4 Sparse Attention Analysis
Spikformer [9] 95.12 82.75 82.83 82.10
SDT [12] 95.77 85.37 83.77 84.30
Spikformer + SEMM [34] 94.98 85.81 83.04 83.37
@@
Spikformer [9] 95.57 93.43 89.97
SDT [12] 96.38 94.68 87.33
Spikformer + SEMM [34] 95.83 93.37 84.95
@@
To test whether stronger attention can compensate for weaker convolutional backbones, we replace SSA with SDSA-v3 [14, 48], where QKV are generated using depthwise separable convolutions:
@@ 5.5 Energy Efficiency Modeling
  • Quantized ANNs are underestimated in efficiency. Bit-serial execution in low-bitwidth ANNs [49, 50] can transform MACs into sequences of ACs, which can exploit bit-level sparsity to skip ineffectual operations—similar to spike sparsity in SNNs. This makes quantized ANNs significantly more efficient than previously assumed.
@@
Table 9: Energy analysis modeling. $F_{Conv}$ and $F_{Mlp}$ denote FLOPs of Conv and MLP modules in ANNs. $B$ is the quantized bit-width in quantized Transformers; $T$ is the time steps in spiking Transformers. $R_{s}$ (firing rate) and $R_{b}$ (bit rate) represent spike sparsity and bit-level sparsity of the quantized activation. $E_{Mac}=4.6pJ$, $E_{Ac}=0.9pJ$, and $E_{Mem}=3.12pJ$ denote energy per MAC, AC, and memory access (per bit energy access from a 16MB cache), respectively [51].
@@
$f$(Q,K,V) Compute $E_{Mac}2N^{2}D$ $BR_{b}\cdot E_{Ac}2N^{2}D$ $TR_{s}\cdot E_{Ac}ND$

### v2

@@ 1 Introduction
Spiking Neural Networks (SNNs) are a biologically inspired paradigm that simulate neural information processing via discrete spikes. These networks excel not only at static image tasks but also in modeling dynamic and temporally structured data [maass1997networks]. Their event-driven nature contributes to high energy efficiency and strong biological plausibility. However, applying SNNs to deep learning architectures—particularly Transformers—remains challenging due to their non-differentiability, limited scalability, and training instability.
@@
In parallel, Artificial Neural Networks (ANNs) have seen tremendous advances through architectural innovations. ResNet [he2016deep] introduced residual learning to ease optimization in deep networks, while Recurrent Neural Networks (RNNs) captured sequential dependencies. The Transformer architecture [vaswani2017attention] unified these advances by leveraging self-attention, enabling efficient parallel modeling of long-range dependencies. Vision Transformer (ViT) [dosovitskiy2020image] further demonstrated the potential of attention mechanisms in visual tasks. Drawing inspiration from these architectures, the SNN community has proposed Spiking ResNet [hu2021spiking], SEW-ResNet [fang2021deep], and spiking RNN variants [xing2020new, xu2024rsnn]. Recently, attention-based spiking models such as Spikformer [zhou2022spikformer], QKFormer [zhou2024qkformer], and SpikingResformer [shi2024spikingresformer] have emerged. The Spike-Driven Transformer series [yao2023spike, yao2024spike, yao2025scaling] improves both efficiency and scalability, enabling applications in image segmentation and object detection.
@@
Despite advancements, several key challenges persist in Spiking Transformers (STs). First, the performance gap between STs and traditional ANNs remains unclear, especially regarding their unique advantages on temporal or relatively complicated data. A systematic evaluation across diverse datasets—static (e.g., ImageNet), event-driven (e.g., DVS-CIFAR10), and sequential (e.g., SCIFAR10)—is essential for assessing their potential. Second, STs consist of multiple interacting components, including spike encoders, neuron models, surrogate gradients, attention modules, and MLP heads, yet the contribution of each module is underexplored. Module-wise ablation is critical for understanding trade-offs and guiding optimization. Third, while SNNs inherently offer energy benefits through sparse, binary spike-based computation, direct comparisons to quantized Transformers are scarce. Quantifying the energy-performance trade-off is necessary to assess the practical utility of spiking models. Moreover, inconsistencies across development frameworks, such as SpikingJelly [fang2023spikingjelly], BrainCog [zeng2023braincog], and BrainPy [wang2023brainpy], further hinder progress by complicating reproducibility, hyperparameter tuning, and fair model comparison. Currently, no unified platform exists for evaluating Spiking Transformers across tasks like classification, segmentation, and detection.
@@
To address these challenges, we introduce the Spiking Transformer Evaluation Platform (STEP), a unified benchmarking framework for building, evaluating, and comparing Spiking Transformers. STEP integrates representative implementations, supports modular component replacement, and enables consistent evaluation across visual tasks. It provides both training-from-scratch and pretraining–finetuning pipelines, and supports integration with backends such as SpikingJelly, BrainCog, and BrainPy. Moreover, leveraging MMSegmentation [mmseg2020] and MMDetection [mmdetection], STEP extends support to dense prediction tasks. Our main contributions are as follows:
@@ 2 Preliminary
To enable spike-based processing, input signals are transformed into temporal spike trains via encoding schemes such as direct, rate, time-to-first-spike (TTFS), and phase encoding [adrian1926impulses, park2020t2fsnn, kim2018deep]. A detailed overview of encoding methods is provided in Appendix A.1.
@@
Spiking neurons transmit information via discrete spikes triggered by membrane potential dynamics. The Leaky Integrate-and-Fire (LIF) model [hunsberger2015spiking] is widely used due to its simplicity and biological plausibility:
[…]
Variants like PLIF [fang2021incorporating] and GLIF [yao2022glif] enhance adaptability with learnable decay or gated mechanisms. Further details are provided in Appendix A.2.
@@
Recent ST models propose lightweight attention [yao2023spike, yao2024spike], hierarchical designs (e.g., QKFormer [zhou2024qkformer]), and multi-task heads (e.g., FCN [long2015fully], FPN [lin2017feature]) to enhance performance across modalities. These improvements drive STs toward practical deployment while retaining neuromorphic efficiency.
@@ 3.2 Broad Dataset Compatibility
Our benchmark supports a wide spectrum of datasets, encompassing static (e.g., ImageNet [deng2009imagenet]), event-based (e.g., DVS-CIFAR10 [li2017cifar10]), and sequential inputs. It also integrates sequential classification tasks to assess temporal modeling capabilities. For dense prediction tasks, we provide plug-and-play support for SNN-adapted segmentation and detection models, such as FCN [long2015fully] and FPN [lin2017feature], based on MMSeg and MMDet toolchains. 3D point cloud and DVS video detection(PKU-DAVIS-SOD [li2023sodformer]) are also supported in the latest version. However, since these datasets are not compatible with most models, we do not conduct unified evaluations on them. Some experimental results can be found in the Appendix B.5.
@@ 3.3 Multi-task Adaptation
While early Spiking Transformers (e.g., Spikformer [zhou2022spikformer], TIM [shen2024tim]) were largely limited to classification, recent efforts such as Spike-Driven Transformer V2 have expanded their scope to include dense vision tasks. Our benchmark extends this trajectory by enabling flexible configuration across classification, segmentation, and detection pipelines within a unified framework.
@@ 3.4 Backend-Agnostic Integration
To enhance accessibility and reusability, our benchmark supports multiple backends such as SpikingJelly [fang2023spikingjelly], BrainCog [zeng2023braincog], and BrainPy. This backend-agnostic design ensures broad compatibility and enables cross-framework reproducibility. Overall, the framework is robust, extensible, and task-agnostic, offering a solid foundation for developing and evaluating Spiking Transformer architectures.
@@ 4.1 Reproduction
Spikformer [zhou2022spikformer] 128 4 400 95.12 (95.51) 77.37 (78.21)
SDT [yao2023spike] 128 4 400 95.77 (95.60) 78.29 (78.40)
QKFormer [zhou2024qkformer] 128 4 400 96.24 (96.18) 79.72 (81.15)
Spikingformer [zhou2023spikingformer] 128 4 400 95.53 (95.81) 79.12 (79.21)
Spikformer + SEMM [zhou2024spikformer] 128 4 400 94.98 (95.78) 77.59 (79.04)
Spiking Wavelet [fang2024spiking] 128 4 400 95.31 (96.10) 76.99 (79.30)
SGLFormer [zhang2024sglformer] 16 4 400 95.88 (96.76) 80.61 (82.26)
SpikingResformer [zhou2023spikingformer]∗∗ 128 4 400 95.69 (97.40) 79.45 (85.98)
Tab. 1 presents our reproduced results on CIFAR-10 and CIFAR-100 [krizhevsky2009learning]. All models are trained using the same optimizer, learning rate, batch size (unless otherwise constrained), training epochs, and random seed. Experiments are conducted on NVIDIA A100 GPUs with 40GB memory.
Overall, our reproduced results are consistent with the original papers. Some models, such as QKFormer, even outperform their reported results, suggesting strong reproducibility. Discrepancies stem mainly from (i) implementation differences, e.g., SpikingResformer originally uses transfer learning, while our setup employs end-to-end training; and (ii) memory limitations, e.g., SGLFormer requires a smaller batch size. To ensure fairness, we avoid dataset- or model-specific tuning and apply a uniform experimental protocol across all baselines. The metrics commonly used to evaluate the framework and reproduction robustness can be found in the Appendix B.4.
@@ 4.2.1 Classification: ImageNet-1K
For ImageNet-1K [deng2009imagenet] we evaluate only Spikformer [zhou2024spikformer] and QKFormer [zhou2024qkformer]: the former is the seminal Spiking-Transformer baseline, while the latter introduces a hierarchical pyramid and currently delivers SOTA accuracy among SNN-based Transformers. Concentrating our limited GPU budget on these two “end-points” lets us cover the full architectural spectrum without incurring the prohibitive cost of training several similar intermediate models. Because ImageNet-1K is orders of magnitude larger and more complex than CIFAR-10/100—and because Spikformer and QKFormer differ greatly in parameter count and memory footprint—forcing a single batch size and epoch schedule would either overflow A100 GPU memory or demand untenable compute. We therefore keep each model’s published regime (QKFormer: 200 epochs × 32/GPU; Spikformer: 300 epochs × 24/GPU), while unifying every other hyper-parameter under a single script; the differing batch sizes and epoch counts are thus an intentional, resource-aware decision rather than an oversight.
@@ 4.2.2 Segmentation: ADE20K
QKFormer [zhou2024qkformer] 256 4 200 73.88 - - -
Spikformer [zhou2023spikingformer] 192 4 300 73.69 69.80 23.51 31.43
Spikformer + SEMM [zhou2024spikformer] - - - - 63.41 13.13 19.76
SDT [yao2023spike] 384 4 300 71.96 63.45 12.08 17.17
For semantic segmentation on ADE20K [zhou2017scene], only SDT [yao2023spike] had been previously evaluated. We conduct a fair comparison by retraining Spikformer  [zhou2024spikformer]and SEMM [zhou2024spiking] without pretraining. Interestingly, both Spikformer variants outperform SDT under identical settings, despite SDT’s original paper reporting strong performance with pretraining. This implies that with appropriate initialization strategies, Spikformer-based models could potentially surpass existing baselines in dense prediction tasks. The segmentation result can be viewed in Fig. 3
@@ 5.1 Neuron Model Evaluation
To investigate the impact of different spiking neuron types on model performance, we replace the default LIF neuron with five widely used variants: PLIF [fang2021incorporating], CLIF [huang2024clif], GLIF [yao2022glif], and KLIF [jiang2023klif]. These models extend the basic LIF neuron [hunsberger2015spiking] by incorporating enhancements such as learnable time constants, gating mechanisms, and surrogate gradient improvements. To quantify the influence of neuron choice, we replace the default LIF cell with four mainstream variants—PLIF [fang2021incorporating], CLIF [huang2024clif], GLIF [yao2022glif], and KLIF [jiang2023klif]. Each variant augments the canonical LIF formulation [hunsberger2015spiking] with additional biological or optimization benefits, ranging from a learnable membrane constant (PLIF) to gated internal states (GLIF) and surrogate-gradient refinements (CLIF, KLIF). Detailed explaination can be found in Appendix A.2.
@@
Spikformer [zhou2022spikformer] 95.12 95.38 95.41 95.85 96.06
SDT [yao2023spike] 95.77 95.49 95.45 95.63 95.91
Spikformer + SEMM [zhou2024spikformer] 94.98 95.44 95.78 95.59 95.66
@@ 5.2 Sequence Modeling
Recent studies [stan2024learning] question the ability of standard Transformers to model long-range temporal dependencies, prompting alternatives like Spiking SSM [shen2025spikingssms]. Datasets such as sCIFAR [chang2017dilatedrecurrentneuralnetworks] and (p)sMNIST [le2015simplewayinitializerecurrent] serialize 2D images into 1D sequences, emphasizing temporal structure. Spiking SSM processes inputs at the pixel level (e.g., 784 steps for a full MNIST image), incurring high computational costs. To adapt Spiking Transformers for serialized inputs, we replace 2D convolutions in the SPS module with 1D convolutions.
@@
FlexTCN [romero2021flexconv] No 99.62 98.63 80.82
SMPConv [kim2023smpconv] No 99.75 99.10 84.86
LMUformer [liu2024lmuformer] No - 98.55 -
ViT [dosovitskiy2020image] No 98.00 97.73 74.95
ViT[dosovitskiy2020image] + SPS ∗∗ No 99.19 98.19 85.62
SpikingSSM [shen2025spikingssms] Yes 99.60 98.40 -
SpikingLMUformer [liu2024lmuformer] Yes - 97.92 -
Spikformer [zhou2022spikformer] Yes 98.84 97.97 84.26
SDT [yao2023spike] Yes 98.77 97.80 82.31
Spikformer + SEMM [zhou2024spikformer] Yes 99.33 98.46 85.61
@@ 5.3 Encoding Schemes
RGB inputs can be converted to spikes through four encoding schemes: direct, phase, rate, and TTFS [adrian1926impulses, park2020t2fsnn, kim2018deep]. As shown in Tab. LABEL:tab:encoding_test, direct encoding—being lossless and repeating the full image at each timestep—aligns with current Spiking Transformers that compute attention independently, yielding the highest Top-1 accuracy. In contrast, the sparser phase, rate, and TTFS encodings reduce spike density and spatial coherence, leading to lower accuracy and emphasizing the need for temporally aware attention or recurrent designs.
@@ 5.4 Sparse Attention Analysis
Spikformer [zhou2022spikformer] 95.12 82.75 82.83 82.10
SDT [yao2023spike] 95.77 85.37 83.77 84.30
Spikformer + SEMM [zhou2024spikformer] 94.98 85.81 83.04 83.37
@@
Spikformer [zhou2022spikformer] 95.57 93.43 89.97
SDT [yao2023spike] 96.38 94.68 87.33
Spikformer + SEMM [zhou2024spikformer] 95.83 93.37 84.95
@@
To test whether stronger attention can compensate for weaker convolutional backbones, we replace SSA with SDSA-v3 [yao2025scaling, chen2018encoder], where QKV are generated using depthwise separable convolutions:
@@ 5.5 Energy Efficiency Modeling
  • Quantized ANNs are underestimated in efficiency. Bit-serial execution in low-bitwidth ANNs [eckert2018neural, zhao2020bitpruner] can transform MACs into sequences of ACs, which can exploit bit-level sparsity to skip ineffectual operations—similar to spike sparsity in SNNs. This makes quantized ANNs significantly more efficient than previously assumed.
@@
Table 9: Energy analysis modeling. $F_{Conv}$ and $F_{Mlp}$ denote FLOPs of Conv and MLP modules in ANNs. $B$ is the quantized bit-width in quantized Transformers; $T$ is the time steps in spiking Transformers. $R_{s}$ (firing rate) and $R_{b}$ (bit rate) represent spike sparsity and bit-level sparsity of the quantized activation. $E_{Mac}=4.6pJ$, $E_{Ac}=0.9pJ$, and $E_{Mem}=3.12pJ$ denote energy per MAC, AC, and memory access (per bit energy access from a 16MB cache), respectively [horowitz20141].
@@
ff(Q,K,V) Compute $E_{Mac}2N^{2}D$ $BR_{b}\cdot E_{Ac}2N^{2}D$ $TR_{s}\cdot E_{Ac}ND$